# Coloring minimal Cayley graphs

### v1

@@ Abstract.
We show that any minimal Cayley graph of a (finitely generated) generalized dihedral or nilpotent group has chromatic number at most 3, while 4 colors are sometimes necessary for soluble groups. On the other hand we construct graphs of unbounded chromatic number that admit a proper edge coloring such that each cycle has some color at least twice. The latter disproves a conjecture of Babai’78 that would have implied that all minimal Cayley graphs have bounded chromatic number – a problem that remains open.
@@ 1. Introduction
Given a group $\Gamma$ and a connection set $C\subseteq\Gamma$ the (undirected, right) Cayley graph $\mathrm{Cay}(\Gamma,C)$ has vertex set $\Gamma$ and $a,b\in\Gamma$ form an edge if $a^{-1}b\in C$. A Cayley graph is minimal (a.k.a irreducible) if $C$ is an inclusion-minimal generating set of $\Gamma$ and in this paper we only consider the case where $C$ is finite. Minimal Cayley graphs appear naturally: a famous and open problem often attributed to Lovász is (equivalent to) whether minimal Cayley graphs are Hamiltonian, see [27, Section 4]. Other areas in which minimal Cayley graphs naturally occur are the genus of a group, see the book [32] or concerning sensitivity [14, Questions 7.7 and 8.2]. Different aspects of these graphs (sometimes of particular groups) and their distinguishing features compared to general Cayley graphs have been considered in [24, 30, 21, 19, 15, 12]. The present work is motivated by a question that has been brought up by Babai [3, 5]:
@@
Does there exist a constant $c$ such that every minimal Cayley graph has chromatic number at most $c$?
Note that assuming minimality here is essential. Indeed, every graph is an induced subgraph of some Cayley graph [6, 16, 4], this is far from being the case for minimal Cayley graphs, which are known to be sparse [3, 31]. However, there also exist (non-minimal) sparse Cayley graphs (of arbitrary girth) that are expanders and hence have unbounded chromatic number [22]. The chromatic number of random Cayley graphs has been studied by Alon [2] and Green [17] as well as some concrete Cayley graphs have been considered recently [9, 10]. However, none of these are minimal. Concerning, the chromatic number of minimal Cayley graphs in [5, Section 3.4] Babai mentions the following:
@@
Hence, this in particular would imply a negative answer to Question 1.1. However, previously in [3] Babai proposed an opposing conjecture, for which we introduce some notions first. A graph $G$ is called no lonely color if $G$ admits an edge coloring satisfying:
@@
Conjecture 1.3.
There is a constant $c$ such that any one popular color graph (no lonely color graph) has chromatic number at most $c$.
Note that Conjecture 1.3 in particular implies that (semi)minimal Cayley graphs have bounded chromatic number and hence would disprove the more recent Conjecture 1.2 and give a positive answer to Question 1.1.
[…]
This is the starting point for the present work: First, we observe that every group admits a minimal Cayley graph of chromatic number at most $3$ and characterize the bipartite case (Theorem 2.2). Then we show that minimal Cayley graphs of finitely generated nilpotent groups (Theorem 2.5) as well as generalized dihedral groups (Theorem 2.6) have chromatic number at most $3$. On the other hand, we show a minimal Cayley graph of a soluble group of chromatic number $4$ and a semiminimal Cayley graph of a nilpotent group of chromatic number $7$ (see Figure 1). Finally, we disprove the stronger statement of Conjecture 1.3 by constructing a family of one popular color graphs of unbounded chromatic number (Theorem 3.2). We further remark, that the clique number of a semiminimal Cayley graph is at most $4$ (Proposition 3.1).
@@ 2. Positive results
Let us begin with the question whether every group admits a minimal Cayley graph of bounded chromatic number. This was already answered by Babai [3, 5]. Indeed, he proved that every group has a 3-colorable minimal Cayley graph contingent upon the truth of some facts about simple groups that were conjectured at that time (see [3, Conjecture 4.3]). Years later the classification of finite simple groups was settled and those conjectures were verified. For completeness we include a proof following Babai’s lines. At the basis of this we have the following [3, Lemma 4.2], which will also use later on. For a group $\Gamma$ denote its minimum chromatic number y $\chi_{\mathrm{min}}(\Gamma)$ the minimum $\chi(\mathrm{Cay}(\Gamma,C))$ over all generating sets $C$ of $\Gamma$.
Lemma 2.1.
@@
Combining this also with ideas present in [12] we get:
Theorem 2.2.
For a group $\Gamma$ we have:
@@
The statement that $\chi_{\mathrm{min}}(\Gamma)=1$ if and only if $\Gamma$ is trivial, is trivial. Now, by Lemma 2.1 $\chi_{\mathrm{min}}(\Gamma)\leq\chi_{\mathrm{min}}(\Gamma/N)$ for any normal subgroup $N$ of $\Gamma$. If $\Gamma$ has a subgroup of index $2$, then it is normal an the quotient is $\mathbb{Z}_{2}$ and has a bipartite Cayley graph. In general, this argument reduces the question to simple groups. Since each such group $\Gamma^{\prime}$ is either cyclic or has a generating set $C^{\prime}$ of size $2$ containing an involution[20], $\mathrm{Cay}(\Gamma^{\prime},C^{\prime})$ has maximum degree $3$ and is different from $K_{4}$. Hence $\chi(\mathrm{Cay}(\Gamma^{\prime},C^{\prime}))\leq 3$. It remains to show that if $\chi_{\mathrm{min}}(\Gamma)=2$, then $\Gamma$ has a subgroup of index $2$. In this case, $\Gamma$ may be partitioned into two independent sets $A,B$, which by vertex transitivity of $\Gamma$ are of equal size. If say $e\in A$, then $A$ consists of all elements of $\Gamma$ that can be expressed as an word of even length in $C$. Hence, $A$ is a subgroup of index $2$. ∎
@@
$\chi(\mathrm{Cay}(\Gamma,C))\leq\max\{\chi(\mathrm{Cay}(\Gamma/\langle C-\{c\}\rangle,c))\mid c\in C\}.$
@@
Since $C=\{c_{1},\ldots,c_{k}\}$ is minimal, for each $c\in C$ the graph $\mathrm{Cay}(\Gamma,C-\{c\})$ is disconnected and its connected components correspond to the vertices of $\mathrm{Cay}(\Gamma/\langle C-\{c\}\rangle,c)$. Moreover, if two vertices $x,y$ of $\mathrm{Cay}(\Gamma,C)$ are connected with an edge corresponding to $c$, then $x,y$ are contained in different components of $\mathrm{Cay}(\Gamma,C-\{c\})$ corresponding to adjacent vertices $x_{c},y_{c}$ of $\mathrm{Cay}(\Gamma/\langle C-\{c\}\rangle,c)$. Hence, if we map every vertex of $x\in\mathrm{Cay}(\Gamma,C)$ to the tuple of vertices $(x_{c_{1}},\ldots,x_{c_{k}})$ we obtain a graph homomorphism into the Cartesian product $\mathrm{Cay}(\Gamma/\langle C-\{c_{1}\}\rangle,c_{1})\square\cdots\square\mathrm{Cay}(\Gamma/\langle C-\{c_{k}\}\rangle,c_{k})$. Hence, the chromatic number of the latter is an upper bound for $\chi(\mathrm{Cay}(\Gamma,C))$. Finally, by a well-known result of Sabidussi [29], the chromatic number of a Cartesian product is the maximum chromatic number of its factors. ∎
@@
Note that Lemma 2.3 alone does not give a useful upper bound for general minimal Cayley graphs. Consider the so-called star graph (see [1],) which is the bipartite graph $\mathrm{Cay}(S_{n},C)$ of the symmetric group of degree $n$ with respect to the set of transpositions involving $1$, i.e., $C=\{(12),(13),\ldots,(1n)\})$. However, $\mathrm{Cay}(S_{n}/\langle C-\{c\}\rangle,c))=K_{n}$ for all $n$ and $c\in C$.
@@
Consider ${\rm Cay}(\Gamma,C)$ a minimal Cayley graph of a Dedekind group $\Gamma$. Since $\Gamma$ is Dedekind, for all $c\in C$ we have that $\langle C-\{c\}\rangle$ is a normal subgroup of $\Gamma$. Hence, $\Gamma/\langle C-\{c\}\rangle$ is a cyclic group generated by the coset of $c$, i.e., the Schreier coset graph (which is a Cayley graph) is a cycle and $\chi({\rm Cay}(\Gamma/\langle C-\{c\}\rangle,c))\leq 3$. The statement follows from Lemma 2.3. ∎
@@
By (b) and Theorem 2.5, it follows that ${\rm Cay}((\Gamma\times\{0\})/H,C_{0})$ is $3$-colorable. We consider $\tilde{f}:(\Gamma\times\{0\})/H\longrightarrow\{0,1,2\}$ a proper coloring. By (a), we have that $\tilde{f}:\Gamma\times\{0\}\longrightarrow\{0,1,2\}$ defined as $f(g,0):=\tilde{f}((g,0)*H)$ is also a proper coloring of ${\rm Cay}(\Gamma\times\{0\},C_{0})$. Now we choose $(y,1)\in C_{1}$ and consider
@@
    If $t_{1}=t_{2}=0$, then $(g_{1},0)$ and $(g_{2},0)$ are adjacent in ${\rm Cay}(\Gamma\times\{0\},C_{0})$ and, hence, $h(g_{1},0)=f(g_{1},0)\neq f(g_{2},0)=h(g_{2},0)$.
@@
For the next result, we denote by $\Phi(\Gamma)$ the Frattini subgroup of $\Gamma$, that is, the intersection of all maximal proper subgroups of $\Gamma$, or $\Phi(\Gamma)=\{e\}$ if it has no maximal subgroups.
@@
For $\Gamma$ a group and $C$ any minimal generating set. The following remarkable properties of the Frattini subgroup are well known (see, e.g., [28, Section 5.2]):
@@
$\Gamma^{\prime}=\{x\cdot y\cdot x^{-1}\cdot y^{-1}\,|\,x,y\in\Gamma\}.$
@@
A group $\Gamma$ satisfies that $\Gamma^{\prime}<\Phi(\Gamma)$ if and only if all its maximal subgroups have prime index (see [25, Theorem A] for other characterizations of these groups). In particular, every nilpotent group satisfies that $\Gamma^{\prime}\leq\Phi(\Gamma)$. A result of Wielandt [28, 5.2.16] proves that for a finite group $\Gamma$, one has that $\Gamma^{\prime}\leq\Phi(\Gamma)$ if and only if $\Gamma$ is nilpotent. However, there are non-nilpotent infinite groups whose commutator is contained in its corresponding Frattini subgroup. A famous such group is the Grigorchuk group, which is a finitely generated $2$-group in which all maximal subgroups have index $2$. In fact $\Gamma^{\prime}=\Phi(\Gamma)$, and has index $8$ in $\Gamma$ [18]. Other examples are considered in [13].
@@
Now suppose $C$ minimal. Again by Lagrange’s Theorem, for any $c\in C$ the subgroup $\langle C-\{c\}\rangle$ has order at least $2^{k-1}$. Hence, the Schreier graph of this subgroup has at most $\frac{n}{2^{k-1}}$ vertices, and $\chi(\mathrm{Cay}(\Gamma/\langle C-\{c\}\rangle,c))\leq\frac{n}{2^{k-1}}$. Thus, with the first part and Lemma 2.3 we have that $\chi(\mathrm{Cay}(\Gamma,C))\leq\min(2k,\frac{n}{2^{k-1}})$, which is maximised exactly if $k=W_{b}(n)$. By definition we have that $W_{b}(n)=\log_{2}(\frac{n}{W_{b}(n)})=\log_{2}(\frac{n}{\log_{2}(\frac{n}{W_{b}(n)})})$. Moreover we clearly have clearly $W_{b}(n)<\log_{2}(n)$, then we finally get that $W_{b}(n)<\log_{2}(\frac{n}{\log_{2}(\frac{n}{\log_{2}(n)})})=\log_{2}n-\log_{2}\log_{2}(\frac{n}{\log_{2}n})$. ∎
@@ 3. Lower bounds
Suppose a one popular color coloring of $K_{6}$ and consider an edge $ab$ of color $1$ and the four triangles $abc,abd,abe,abf$. At most two among these four triangles have popular color $1$. If two of these triangles, say $abc,abd$ have one popular color among different colors $2,3$, then the triangle $acd$ has no popular color. Hence two triangles say $abc,abd$ have popular color $2$ then the triangles $abe,abf$ must have popular color $1$. So assume that the edge $ae$ is colored in $1$. But then the edges $ec$ and $ed$ both must be of color $1$ to make triangles $aec$ and $aed$ have a popular color. But then the degree of $e$ in color $1$ is $3$.
@@
We finish disproving [3, Conjecture 3.5], i.e., that one popular color graphs have bounded chromatic number. The construction we provide is based on one of the fundamental constructions for triangle-free graphs of arbitrary chromatic number due to Tutte (alias Blanche Descartes) and independently Zykov, see Nešetřil’s survey [26] for a more detailed discussion.
@@
We proceed by induction on $k\geq 1$. Chose $G_{1}$ to be the graph with a single vertex and $G_{2}=C_{4}$ with edges colored alternatingly with two colors. If $k\geq 3$ then denote by $n$ the order of $G_{k-1}$. Define $X$ as a set of size $(k-1)(n-1)+1$. Now, for every subset of $Y$ size $n$ of $X$ take a copy $G^{\prime}$ of $G_{k-1}$ (where all copies can be considered to be edge-colored with the same set) and add a perfect matching between $Y$ and the copy $G^{\prime}$. Each of these new matchings will be edge colored with its own private color.
@@
The fact that $\chi(G_{k})\geq k$ is well-known, see e.g. [26].
While we have disproved the strong variant of Conjecture 1.3 its weak variant remains open. Let us propose a strengthening of it.
@@
Note that this conjecture for $d=2$ is the weak variant of Conjecture 1.3. However, it is open even for the case $d=1$.

### v2

@@ Abstract.
In 1978 Babai raised the question whether all minimal Cayley graphs have bounded chromatic number; in 1994 he conjectured a negative answer. In this paper we show that any minimal Cayley graph of a (finitely generated) generalized dihedral or nilpotent group has chromatic number at most 3, while 4 colors are sometimes necessary for soluble groups. On the other hand we address a related question proposed by Babai in 1978 by constructing graphs of unbounded chromatic number that admit a proper edge coloring such that each cycle has some color at least twice. The latter can be viewed as a step towards confirming Babai’s 1994 conjecture – a problem that remains open.
@@ 1. Introduction
Given a group $\Gamma$ and a connection set $C\subseteq\Gamma$ the (undirected, right) Cayley graph $\mathrm{Cay}(\Gamma,C)$ has vertex set $\Gamma$ and $a,b\in\Gamma$ form an edge if $a^{-1}b\in C$. Note that we will consider Cayley graphs as undirected even if the connection set is not inverse-closed. A Cayley graph is minimal (a.k.a irreducible) if $C$ is an inclusion-minimal generating set of $\Gamma$ and in this paper we only consider the case where $C$ is finite. Minimal Cayley graphs appear naturally: a famous and open problem often attributed to Lovász is (equivalent to) whether minimal Cayley graphs are Hamiltonian, see [28, Section 4]. Other areas in which minimal Cayley graphs naturally occur are the genus of a group, see the book [33] or concerning sensitivity [15, Questions 7.7 and 8.2]. Different aspects of these graphs (sometimes of particular groups) and their distinguishing features compared to general Cayley graphs have been considered in [25, 31, 22, 20, 16, 13]. The present work is motivated by a question that has been brought up by Babai [3, 5]:
@@
Does there exist a finite constant $c$ such that every minimal Cayley graph has chromatic number at most $c$?
Babai conjectured in 1994 a negative answer even in the case the group is finite. Note that assuming minimality here is essential. Indeed, every graph is an induced subgraph of some Cayley graph [6, 17, 4], this is far from being the case for minimal Cayley graphs, which are known to be sparse [3, 32]. However, there also exist (non-minimal) sparse Cayley graphs (of arbitrary girth) that are expanders and hence have unbounded chromatic number [23]. The chromatic number of random Cayley graphs has been studied by Alon [2] and Green [18] as well as some concrete Cayley graphs have been considered recently [9, 10]. However, none of these are minimal. Concerning the chromatic number of minimal Cayley graphs, in [5, Section 3.4] Babai mentions the following:
@@
Hence, this in particular would imply a negative answer to Question 1.1. However, previously in [3] Babai had proposed an opposing conjecture (which he withdrew in 1994 when he proposed Conjecture 3.4), for which we introduce some notions first. A graph $G$ is called no lonely color if $G$ admits an edge coloring satisfying:
@@
Question 1.3.
Is there a finite constant $c$ such that any one popular color graph (no lonely color graph) has chromatic number at most $c$?
Note that a positive answer to Question 1.3 would imply that (semi)minimal Cayley graphs have bounded chromatic number, and hence would disprove the more recent Conjecture 3.4 and give a positive answer to Question 1.1.
This is the starting point for the present work: First, we observe that every group admits a minimal Cayley graph of chromatic number at most $3$ and characterize the bipartite case (Theorem 2.2). Then we show that minimal Cayley graphs of finitely generated nilpotent groups (Theorem 2.5) as well as generalized dihedral groups (Theorem 2.6) have chromatic number at most $3$. On the other hand, we show a minimal Cayley graph of a soluble group of chromatic number $4$ and a semiminimal Cayley graph of a nilpotent group of chromatic number $7$ (see Proposition 3.1). Finally, we negatively answer the stronger statement of Question 1.3 by constructing a family of one popular color graphs of unbounded chromatic number (Theorem 3.3). We further remark, that the clique number of a semiminimal Cayley graph is at most $4$ (Proposition 3.2).
@@ 2. Positive results
Let us begin with the question whether every finite group admits a minimal Cayley graph of bounded chromatic number. This was already answered by Babai [3, 5]. Indeed, he proved that every finite group has a 3-colorable minimal Cayley graph. His proof depends on the fact that every finite simple group can be generated by two elements, one of which is an involution. This is a known consequence of the Classification of Finite Simple Groups. For completeness we include a proof following Babai’s lines. At the basis of this we have the following [3, Lemma 4.2], which will also use later on. For a group $\Gamma$ denote its minimum chromatic number and $\chi_{\mathrm{min}}(\Gamma)$ the minimum $\chi(\mathrm{Cay}(\Gamma,C))$ over all generating sets $C$ of $\Gamma$.
Lemma 2.1 ([4, Lemma 4.1]).
@@
Using this we get:
Theorem 2.2 (Babai).
For a finite group $\Gamma$ we have:
@@
The statement that $\chi_{\mathrm{min}}(\Gamma)=1$ if and only if $\Gamma$ is trivial, is trivial. Now, by Lemma 2.1 $\chi_{\mathrm{min}}(\Gamma)\leq\chi_{\mathrm{min}}(\Gamma/N)$ for any normal subgroup $N$ of $\Gamma$. If $\Gamma$ has a subgroup of index $2$, then it is normal and the quotient is $\mathbb{Z}_{2}$ and has a bipartite Cayley graph. In general, this argument reduces the question to simple groups. Since each such group $\Gamma^{\prime}$ is either cyclic or has a generating set $C^{\prime}$ of size $2$ containing an involution[21], $\mathrm{Cay}(\Gamma^{\prime},C^{\prime})$ has maximum degree $3$ and is different from $K_{4}$. Hence $\chi(\mathrm{Cay}(\Gamma^{\prime},C^{\prime}))\leq 3$. It remains to show that if $\chi_{\mathrm{min}}(\Gamma)=2$, then $\Gamma$ has a subgroup of index $2$. In this case, $\Gamma$ may be partitioned into two independent sets $A,B$, which by vertex transitivity of $\Gamma$ are of equal size. If say $e\in A$, then $A$ consists of all elements of $\Gamma$ that can be expressed as a word of even length in $C$. Hence, $A$ is a subgroup of index $2$. ∎
@@
$\chi(\mathrm{Cay}(\Gamma,C))\leq\max\{\chi(\mathrm{Cay}(\Gamma/\langle C-\{c\}\rangle,\{c\}))\mid c\in C\}.$
@@
Since $C=\{c_{1},\ldots,c_{k}\}$ is minimal, for each $c\in C$ the graph $\mathrm{Cay}(\Gamma,C-\{c\})$ is disconnected and its connected components correspond to the vertices of $\mathrm{Cay}(\Gamma/\langle C-\{c\}\rangle,\{c\})$. Moreover, if two vertices $x,y$ of $\mathrm{Cay}(\Gamma,C)$ are connected with an edge corresponding to $c$, then $x,y$ are contained in different components of $\mathrm{Cay}(\Gamma,C-\{c\})$ corresponding to adjacent vertices $x_{c},y_{c}$ of $\mathrm{Cay}(\Gamma/\langle C-\{c\}\rangle,\{c\})$. Hence, if we map every vertex of $x\in\mathrm{Cay}(\Gamma,C)$ to the tuple of vertices $(x_{c_{1}},\ldots,x_{c_{k}})$ we obtain a graph homomorphism into the Cartesian product $\mathrm{Cay}(\Gamma/\langle C-\{c_{1}\}\rangle,\{c_{1}\})\square\cdots\square\mathrm{Cay}(\Gamma/\langle C-\{c_{k}\}\rangle,\{c_{k}\})$. Hence, the chromatic number of the latter is an upper bound for $\chi(\mathrm{Cay}(\Gamma,C))$. Finally, by a well-known result of Sabidussi [30], the chromatic number of a Cartesian product is the maximum chromatic number of its factors. ∎
@@
Note that Lemma 2.3 alone does not give a useful upper bound for general minimal Cayley graphs. Consider the so-called star graph (see [1]), which is the bipartite graph $\mathrm{Cay}(S_{n},C)$ of the symmetric group of degree $n$ with respect to the set of transpositions involving $1$, i.e., $C=\{(12),(13),\ldots,(1n)\})$. However, $\mathrm{Cay}(S_{n}/\langle C-\{c\}\rangle,\{c\}))=K_{n}$ for all $n$ and $c\in C$.
@@
Consider ${\rm Cay}(\Gamma,C)$ a minimal Cayley graph of a Dedekind group $\Gamma$. Since $\Gamma$ is Dedekind, for all $c\in C$ we have that $\langle C-\{c\}\rangle$ is a normal subgroup of $\Gamma$. Hence, $\Gamma/\langle C-\{c\}\rangle$ is a cyclic group generated by the coset of $c$, i.e., the Schreier coset graph (which is a Cayley graph) is a cycle and $\chi({\rm Cay}(\Gamma/\langle C-\{c\}\rangle,\{c\}))\leq 3$. The statement follows from Lemma 2.3. ∎
@@
By (b) and Theorem 2.5, it follows that ${\rm Cay}((\Gamma\times\{0\})/H,C_{0})$ is $3$-colorable. We consider $\tilde{f}:(\Gamma\times\{0\})/H\longrightarrow\{0,1,2\}$ a proper coloring. By (a), we have that ${f}:\Gamma\times\{0\}\longrightarrow\{0,1,2\}$ defined as $f(g,0):=\tilde{f}((g,0)*H)$ is also a proper coloring of ${\rm Cay}(\Gamma\times\{0\},C_{0})$. Now we choose $(y,1)\in C_{1}$ and consider
@@
    If $t_{1}=t_{2}=0$, then $(g_{1},0)$ and $(g_{2},0)$ are adjacent in ${\rm Cay}(\Gamma\times\{0\},C_{0})$ and hence, $h(g_{1},0)=f(g_{1},0)\neq f(g_{2},0)=h(g_{2},0)$.
@@
This concludes the proof. ∎
For the next result, we denote by $\Phi(\Gamma)$ the Frattini subgroup of $\Gamma$, that is, the intersection of all maximal proper subgroups of $\Gamma$, or $\Phi(\Gamma)=\{e\}$ if it has no maximal proper subgroups.
@@
For $\Gamma$ a group and $C$ any minimal generating set. The following remarkable properties of the Frattini subgroup are well known (see, e.g., [29, Section 5.2]):
@@
$\Gamma^{\prime}=\langle x\cdot y\cdot x^{-1}\cdot y^{-1}\,|\,x,y\in\Gamma\rangle.$
@@
A group $\Gamma$ satisfies that $\Gamma^{\prime}<\Phi(\Gamma)$ if and only if all its maximal subgroups have prime index (see [26, Theorem A] for other characterizations of these groups). In particular, every nilpotent group satisfies that $\Gamma^{\prime}\leq\Phi(\Gamma)$. A result of Wielandt [29, 5.2.16] proves that for a finite group $\Gamma$, one has that $\Gamma^{\prime}\leq\Phi(\Gamma)$ if and only if $\Gamma$ is nilpotent. However, there are non-nilpotent infinite groups whose commutator is contained in its corresponding Frattini subgroup. A famous such group is the Grigorchuk group, which is a finitely generated $2$-group in which all maximal subgroups have index $2$. In fact $\Gamma^{\prime}=\Phi(\Gamma)$, and has index $8$ in $\Gamma$ [19]. Other examples are considered in [14].
@@
Now suppose $C$ minimal. Again by Lagrange’s Theorem, for any $c\in C$ the subgroup $\langle C-\{c\}\rangle$ has order at least $2^{k-1}$. Hence, the Schreier graph of this subgroup has at most $\frac{n}{2^{k-1}}$ vertices, and $\chi(\mathrm{Cay}(\Gamma/\langle C-\{c\}\rangle,\{c\}))\leq\frac{n}{2^{k-1}}$. Thus, with the first part and Lemma 2.3 we have that $\chi(\mathrm{Cay}(\Gamma,C))\leq\min(2k,\frac{n}{2^{k-1}})$, which is maximised exactly if $k=W_{b}(n)$. By definition we have that
$W_{b}(n)=\log_{2}(\frac{n}{W_{b}(n)})=\log_{2}(\frac{n}{\log_{2}(\frac{n}{W_{b}(n)})}).$
Moreover, we clearly have clearly $W_{b}(n)<\log_{2}(n)$, then we finally get that
$W_{b}(n)<\log_{2}(\frac{n}{\log_{2}(\frac{n}{\log_{2}(n)})})=\log_{2}n-\log_{2}\log_{2}(\frac{n}{\log_{2}n}),$
which concludes the proof. ∎
@@ 3. Lower bounds
We begin this section by presenting a minimal Cayley graph of chromatic number $4$ and a semiminimal Cayley graphs of chromatic number $7$.
Proposition 3.1.
There exist minimal Cayley graphs of chromatic number $4$ and semiminimal Cayley graphs of chromatic number $7$.
Proof.
[…]
Figure 1. The graphs $\mathrm{Cay}(\mathbb{Z}_{3}\rtimes\mathbb{Z}_{7},\{(0,1),(1,0)\})$ and $G$ such that $G\boxtimes K_{2}=\mathrm{Cay}(Q_{32},C)$ from Proposition 3.1.
@@
Suppose that there exists a one popular color coloring of $K_{6}$ and consider an edge $ab$ of color $1$ and the four triangles $abc,abd,abe,abf$. At most two among these four triangles have popular color $1$. If two of these triangles, say $abc,abd$ have one popular color among different colors $2,3$, then the triangle $acd$ has no popular color. Hence two triangles say $abc,abd$ have popular color $2$ then the triangles $abe,abf$ must have popular color $1$. So assume that the edge $ae$ is colored in $1$. But then the edges $ec$ and $ed$ both must be of color $1$ to make triangles $aec$ and $aed$ have a popular color. But then the degree of $e$ in color $1$ is $3$.
@@
We finish by giving a negative answer to the stronger version of Question 1.3, i.e., we provide a family of one popular color graphs with unbounded chromatic number. The construction we propose is based on one of the fundamental constructions for triangle-free graphs of arbitrary chromatic number due to Tutte (alias Blanche Descartes) [12], also see Nešetřil’s survey [27] for a more detailed discussion.
@@
We proceed by induction on $k\geq 1$. Chose $G_{1}$ to be the graph with a single vertex and $G_{2}=C_{4}$ with edges colored alternatingly with two colors. If $k\geq 3$ then denote by $n$ the number of vertices of $G_{k-1}$. Define $X$ as a set of $(k-1)(n-1)+1$ new vertices without edges. Now, for every subset of $Y$ size $n$ of $X$ take a copy $G^{\prime}$ of $G_{k-1}$ (where all copies can be considered to be edge-colored with the same set) and add a matching of size $n$ that connects the vertices of $Y$ with the vertices of $G^{\prime}$. Each of these new matchings will be edge colored with its own private color. The resulting graph is $G_{k}$.
@@
The fact that $\chi(G_{k})\geq k$ is well-known, see e.g. [12, 27]. ∎
While we have negatively answered the one-popular-color variant of Question 1.3 its no-lonely-color variant remains open. Let us propose a strengthening of it.
@@
Note that this conjecture for $d=2$ implies a positive answer to the weak variant of Question 1.3. However, it is open even for the case $d=1$. On the other hand, it also remains open if graphs with a one-popular-color coloring exist that satisfy Conjecture . Finally, note that the questions about the boundedness of the chromatic number of minimal or semiminimal Cayley graphs of general finite groups remain open.